

********************************************************************
UARS WEEKLY STATUS REPORT - 08 JAN 1999
********************************************************************
***********
SOLSTICE
***********
SOLSTICE continues to operate normally, collecting solar data on all
available orbital tracking periods and stellar calibration observations,
whenever possible. The CDHF has completed SOLSTICE version 10 re-processing
and is now running the current processing stream only.
***********
SUSIM
***********
Nothing received.
***********
PEM
***********
PEM is now under PMON control! OBC activation tests of the high energy
particle spectrometer (HEPS -- particles greater than 30 keV) were declared
successful on UARS day 2670 (January 2, 1999). Beginning this date, all PEM
activations should yield correct scientific data; however, data generated from
the HEPS subsystem between December 17, 1998 and January 1, 1999 should be
considered invalid and not used. The PEM team would like to thank the FOT, and
in particular Peter Thomas, for their work and efforts related to obtaining a
correct OBC activation procedure for the HEPS subsystem.
***********
HALOE
***********
HALOE was powered off on January 6th for the upcoming yaw after
successfully operating since December 17, 1998. All engineering parameters
were nominal and the science data products continue to be of excellent quality.
***********
MLS
***********
The instrument is now off as scheduled.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
Routine processing of instrument Daily Activity Plans continues
with no problems. The MPG continues to assist the Project
Scientist with the scheduling of instrument operations.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 26 December 1998 (Orbit 39838, GMT Day 360) through 01
January 1999.
The observatory is now in Forward Flight (Southern Hemisphere viewing) with
a +0.5 degree Roll (cold side down) and is performing nominally under two-
battery operations. The instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
MLS 363/2312 39896 Off
ACRIM 363/2312 39896 On
SUSIM 364/0015 39897 On
PEM 365/1626 39922 Testing
A summary of Instrument operations supplied by the Mission Planning Group
for the month December 1998 is contained in Appendix F, UARS Instrument
Operations Summary.
The beta angle ranged between 32.2 degrees to 29.9 degrees for this report
period. The beta angle is now increasing toward a minimum angle of +1.0
degrees on 11 January 1999 (DOY 011).
UARS Flight Software Time was adjusted consistent with the United States
Naval Observatory (USNO) Time Service announcement advising that the
International Earth Rotation Service (IERS) will retard UTC time by 1.0
second at the end of December 1998. This meant the addition of one leap
second as the last second of 1998.
The PEM Instrument automated turn-on procedure for Day-Only Operation has
been completed and successfully tested in orbit. HEPS-1 and HEPS-2 are
working (HEPS-3 will not be operated day only).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.8 V end-of-night (EON) load bus voltage.
Battery temperatures are stable with a temperature delta between Battery 2
and Battery 3 of 1.0 to 1.3 degrees C. Additional battery performance data
for each battery on 25 December 1998 and 01 January 1999 are provided in
Appendix D.
The clock error ranged from 10.3 to -11.4 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT for
this report. A summary is listed below and the details of ALERT and Out-
Of-Limits occurrences are listed in Appendix B.
% ALERTS = 05
% OUT-OF-LIMITS = 02
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
% FOT 05
% ACRIM 01
% MLS 03
% PEM 09
% SOLSTICE 03
% SUSIM 01
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
% FOT 01
Three (3) Anomaly Report remained OPEN at the end of this report period
(see Appendix E).
TRANSPONDER FREQUENCIES:
Transponder A: 98/286 - 2287.496936
Transponder B: 98/300 - 2287.498964
GROUND SYSTEM OPERATION
One (1) operational support problems occurred this period. Anomaly Reports
are included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from
mission start through UTC Day 98/365 show the observatory data loss to be
21 hours, 40 minutes, 58.964 seconds (an increase of 30.432 seconds since
last report period). This is a 0.0340 percent data loss which equals a
99.9660 percent data capture for the mission.
The increase of 30.432 seconds occurred over two (2) days:
7 0.512 seconds on Day 362 due to data hit; and
7 29.920 seconds on Day 365 due to data hit, year change, leap second
adjust.
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
The SSPP temperature stability is being analyzed. It appears that since
SOLSTICE is cycling, the usual amount of heat is not being generated and
along with the other SSPP instruments, getting slightly colder.
Investigation revealed a problem in operation of Compensation Heater 18B.
Operational changes were implemented but sufficient operational time to
evaluate the thermal effect has not elapsed.
The FOT is investigating the feasibility of a new PMON action to change
battery V/T curves based on loading to mitigate battery risks in sun
pointing safehold.
YEAR 2000 OPERABILITY SUPPORT - No change
UARS PROPULSION MODULE B/D THRUSTER LINE FILL
Planning continues for the simulator testing.
UARS RE-ENGINEERING - TELEMETRY PROCESSING - No change
MEETINGS
None
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw Around Maneuver is scheduled for scheduled for 11 January
1999.
The next annular solar eclipse will occur on 16 February 1999 over the
Indian Ocean, Australia and the South Pacific Ocean.
PERSONNEL CHANGES
none
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
Appendix F, UARS Instrument Operations Summary
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 01 January 1999
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
39846 360/1416 98-516 MLS Scan Act OFF
39851 360/2251 98-517 MLS Scan Act ON
39872 362/0740 98-518 OBC Clk Adj fr 0.00 to 20.6 msec/day
39878 362/1757 98-519 PSU Enabled Comp Htr 18 A&B sides
39895 363/2039 98-520 SOLSTICE SOLSTICE RTSs
39896 363/2312 98-521 MLS Power OFF
39896 363/2312 98-522 ACRIM Power ON
39897 364/0015 98-523 SUSIM Power ON
39900 364/0420 98-524 OBC Clk Adj fr 20.6 to 0.00 msec/day
39900 364/0518 98-525 PSU Comp HTR 18 ON BOTH
39901 364/0707 98-526 SOLSTICE SOLSTICE RESET
39903 364/0932 98-526 SOLSTICE SOLSTICE RESET
39907 364/1550 98-537 PSU Comp HTR 18 ON BOTH
39921 365/1403 98-528 PEM Revised DAY-ONLY sequences
39922 365/1626 98-529 PEM DAY-ONLY TEST-1
39923 365/1735 98-530 PEM DAY-ONLY TEST-2
39924 365/1949 98-531 PEM DAY-ONLY TEST-3
39925 365/2126 98-532 PEM DAY-ONLY TEST-4
39927 001/0025 99-001 PEM PMON/PEM Test 6
39928 001/0206 99-002 PEM PMON/PEM Test 7
39929 001/0346 99-003 PEM PMON/PEM Test 8
39930 001/0528 99-004 PEM PMON/PEM Test 9
_______________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 01 January 1999
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
QTSISAIF Y-LO SSPP Baseplate by SISA I/F 39845-912
QTACRIMIF Y-LO SSPP Temp at ACRIM I/F 39845-96
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
ACTEMPHMREDL ACRIM RED LOW 39897-8
ACTEMANA1RDL ACRIM RED LOW 39897-8, 900
ACTEMANA2RDL ACRIM RED LOW 39897-8
SOPUPERSTATA SOLSTICE DETECTORS OFF 39900
SOPUPERSTATA MODE SELECTED= "TWO" RED 39903
_______________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 01 January 1999
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
39872 362/07:47 (28 Dec 98) from 00.0 to 20.6 msec/day
39900 364/04:23 (30 Dec 98) from 20.6 to 00.0 msec/day
_______________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 01 January 1999
25 Dec 1998 - GMT Day 359 Beta = 32.2 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 +61.6/+39.2 -2.73/-3.99 +0.0/ +0.0 16.8 0.000 0.0
2 +11.2/-67.2 +6.11/+5.15 +30.4/-15.2 25.3 1.028 21.8
3 +134.4/11.2 +5.15/+3.87 +30.0/-16.4 25.1 1.024 21.7
01 Jan 1999 - GMT Day 001 Beta = 29.9 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 +56.0/+33.6 -2.73/-3.99 +0.0/ +0.0 16.8 0.000 0.0
2 +33.6/-22.4 +6.44/+5.15 +31.2/-15.2 25.1 1.026 21.8
3 +44.8/-5.6 +5.15/+3.87 +30.4/-16.0 25.1 1.025 21.6
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- -------- ----------
no changes V/T 5
APPENDIX E
UARS ANOMALY REPORTS
Week ending 01 January 1999
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
98-124 FOT 39878 362/1757 20485 No ODM/GCMR capability
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
None this report period
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------- ------ ----- ---- --- -----------------
None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
98-117 FOT Dec 98 353/07:02 14170 Generic Late Acq
98-123 FOT 39788 356/17:45 AP/ODN PROBLEM
98-124 FOT 39878 362/18:00 20485 N0 ODM/GCMR
_______________________________________________________________
APPENDIX F
UARS INSTRUMENT OPERATIONS SUMMARY
Week ending 01 January 1999
INSTRUMENT SUMMARY FOR MONTH OF DECEMBER 1998 (hours)
UNPLANNED UNSCHEDULED
ON Day- OFF OFF OFF
Only (Insufficient (Other
Instrument (Actual) Ops (Scheduled) Power) Problem) COMMENTS
---------- ------ --- ----------- ----------- ----------- --------------
ACRIM 599 145 0 0
HALOE 575 169 0 0
HRDI 216 528 0 0 0
MLS 141 603 0 0
PEM (axis) 142 0 602 0 0
PEM (zeps) 4 715 25 0 0
PEM (heps) 4 0 740 0 0 See Note 1
SOLSTICE 600 144 0 0 0
SUSIM 597 147 0 0
WINDII 235 509 0 0
Note 1. PEM High Energetic Particle System (HEPS) was OFF while
developing and testing automated turn-on procedure. The automated procedure
will operate while the spacecraft is in daylight and has power available
from the solar array. (When spacecraft is in darkness, operation of
instruments and
spacecraft systems rely on power from batteries. Only two of the three
spacecraft batteries are currently operational. Battery power is not
sufficient to operate all instruments simultaneously when spacecraft is in
darkness.).
Cold side of spacecraft was facing northward: 12.5 days
Cold side of spacecraft was facing southward: 18.5 days
SEVNTFW99/01